You are here: Foswiki>Commissioning Web>Logs>25Jan013_Log (29 Jan 2013, JuanGuerra)Edit Attach

UT 20130124 AO#1

Observer:Esposito,S(Arcetri),Bussoni,L(Arcetri),Puglisi,A(Arcetri),Pinna,E(Arcetri),JCGuerra(LBTO),JChristou(LBTO),DMiller(LBTO)
Telescope Operator: Steve..(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JCGuerra(LBTO),JChristou(LBTO),DMiller(LBTO))
Telescope Support: Kevin N(LBTO)
Instrument Support: Tim S (LBTO)

Summary

AGW#2 is in the laboratory. A full test of the wavefront sensor including the CCD47 is done without any problem at laboratory ambient temperature (2 hours) The AGW#1 is down to repair the pic board, study the possible vignetting of the W beam splitter (cube) in the Lucifer FOV and the test of the failure of the stage Y.

Detail

The work in the AGW#2 is centered in making the system working and to see if we don't have glycol loose after the replacement of one of the hose (quick connectors).

The AGW#2 is running for two hours and we don't see any dripping from the connector and also the Jesper guider was repaired one of the connector that it was found broken.

It seems that the connectors was mashed during any handling of the AGW. The hose connector was replaced and sokem protection was put in place to avoid any touch of the connectors.

AGW#1 shows a problem with the stage in Z instead of the Y. A more deep work has to be done and we are focused to have the AGW#2 working for the cool test and to put back in the telescope the day after.

At the end of the day the AGW#2 it is working properly without any failure and it will be ready to be installed tomorrow. -- JuanGuerra - 26 Jan 2013
Topic revision: r2 - 29 Jan 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback